Exhibit 10-9

Marginal Propensity to Consume (MPC)

Marginal Propensity to Save (MPS)

  
Multiplier (m)

0.92

(A)

(B)

(C)

(D)

10

0.85

(E)

(F)

(G)

0.20

(H)

Refer to Exhibit 10-9. What is the value of the marginal propensity to save (MPS) that would correctly fill in blank (E) and the multiplier that would correctly fill in blank (F)?

0.012; 0.83

0.12; 88

0.15; 15

0.15; 6.67
Answer #1

Marginal Propensity to save or MPS = 1 - Marginal Propensity to Consumer (MPC)

For MPC= 0.85, MPS = 1 - 0.85 = 0.15

Thus, the MPS in column (E) will be 0.15.

And Multiplier (m) = 1/1-MPC = 1/MPS = 1/0.15 = 6.67

Thus, m in column (F) will be equal to 6.67.

Therefore, option 4. 0.15; 6.67 is correct.